Night 1 of AZ science
In preparation for it, ran up the mods GUI as LBTO and ran a WarmStart
Warm reset of the grating tilts
>> bgrtilt1 reset
** ERROR: BGRTILT1 BGRTILT1=19885 Invalid RESET for Grating Tilt, Requires a HARDRESET or SOFTRESET
Same with rgrtilt1 reset
>> rgrtilt1 reset
** ERROR: RGRTILT1 RGRTILT1=20930 Invalid RESET for Grating Tilt, Requires a HARDRESET or SOFTRESET
** Abort, Retry or Ignore? > r
Need to change modsWake.pro commands:
- bgrtilt1 reset to bgrtilt1 softreset
- rgrtilt1 reset to rgrtilt1 softreset
opening up - some lightning to West and South.
MODS in observing mode.
BD28 120 sec standard star through 5" slit. dual grating. 120 sec. (did a red-only one too, but not needed for this program).
acqMODS mods.14.J02121.acq (no... went to elevation 1.0!! wrong script to run).
- mselect id531945 (worked - we see mask selected but out... perhaps out from next step...)
script failed with Error: Instconfig red imaging (timed out).
- take 30 sec image through r_sdss with mask out (we think!) YES, mask is out.
- also error -
"failed to receive full guide cam readout wi" (this was reported in IT #3252).
- modsCmd slitmask reset successful
- acqMODS -n Instrument mods.22.J21211.acq worked...
called Torsten ... rebooted a machine (mods1-azcamg). Worked after this.
mask image is 8 and field image 9: confirmatory image is 10.
execMODS mods.22.J21211.obs started!!! Will be 5 20-minute exposures.
- guider telling us 0.9-1.0 seeing. So it should be better on the red images, about 0.5". FWHM ~ 4.3 on field image, number 9.
- seeing is reported as 1.5" on GCS now... worsening. Wind picking up. This is during exposure 3 of 5. Now 1.8"...
- 22:28 lost guiding for a few seconds...
- * with 362 remaining, we stopped the last integration. We were losing guiding too frequently.
Slewing to c04_131_517 (a poor weather target). 1.2" slit, dual grating
- no star found... (was a 13th mag guide star).
- modified acq script to use 11th mag guide star.
- 16 is slit, 17 is field image. 90 seconds. 18 is field image with 180 seconds.
- while ID'ing field, Geno was outside and felt rain drops... moved MODS to calibration mode (closed hatch) and Geno is closing. Wind blowing drops of rain from East. We were pointing to the SSW az =197 degrees and vent doors were shut. RH ~ 46%.
- bias8K ccdbin 1 2 (binned biases). b: 9-18 r: 19-28 (telescope not at zenith)
- flats with MOS mask: boxes saturate... even though grmosflats template was used.
- arcs: some bright lines saturate.
- slitless flats: dual: b: Vflat 5, clear 30K and r: Vflat 5 ND1.5 35K
- 0.3" flats - many are close to saturation. made grflats_03_bin12CORR.cal and cntrl-c megascript.
- ./Sept30rest.multcal will do the 0.3" calibs plus the biases. (quartz lamp flats saturated anyway even with reduced exp time)
modsCmd PARTNER AZ
modsCmd PROPID CALIBRATION
modsCmd PI_NAME CALIBRATION
# 5" x 60 fat slit for standard, dual and red-only (red-only not used)
# ID531945 dual
# grating flats through slits
execMODS -u J2121_flats.cal
execMODS -u J2121_lamps.cal
# slitless grating flats
execMODS -u grslitflat.cal
# imaging flats
# execMODS -u blueimflat.cal
# execMODS -u redimflat.cal
# flats for B. Green program (ccdbin 1 2 and 0.3" slit)
execMODS -u grflats_03_bin12.cal (these are saturated in brightest region of blue and red with QTH1, QTH1+QTH2 hard sat in brightest area of blue with 12 sec, close to sat in red 2 sec).
execMODS -u grslitflat_bin12.cal
execMODS -u grlamps_03_bin12.cal
execMODS -u /home/MODSeng/modsScripts/Support/modsSleep.pro
#execMODS -u bias1K.cal
execMODS -u bias3K.cal
execMODS -u bias8K.cal
- MODS: almost 1 MOS set of F Bian done, but cut short by clouds. Only about 2 hours on sky.
- Worked well except two glitches:
- Glitch at end of night when running biases. One bias script ran, but subsequent one complained at "calmode". Guide probe warning AGw X Y values were in orange.
- Glitch at beginning which must have occurred between setting mask and filter. Set filter manually, but then we had the problem with the incomplete guider readout which required the MODS1-AZCAMg computer to be rebooted.
- 29 Sep 2011